OpenStudy (anonymous):

A multiple choice test consists of eight questions, each of which has five choices. Each question has exactly one correct answer. William guesses randomly at each answer. What is the probability that he gets six or fewer questions correct?

ganeshie8 (ganeshie8):

\(\large P(X\le 6) = 1 - [P(7) + P(8)] \)

ganeshie8 (ganeshie8):

its a binomial distribution, u knw how to find P(7) and P(8) ?

ganeshie8 (ganeshie8):

\(P(7) = ~^8C_7 (\frac{1}{5})^7(\frac{4}{5})\) \(P(8) = (\frac{1}{5})^8\)

ganeshie8 (ganeshie8):

evaluate and plugin